Perdon Corporation manufactures safes—large mobile safes, and large walk-in stationary bank safes. As part of its annual budgeting process, Perdon is analyzing the profitability of its two products. Part of this analysis involves estimating the amount of overhead to be allocated to each product line. The information shown below relates to overhead.

Mobile Safes

Walk-in Safes

Units planned for production 190 40
Material moves per product line 290 190
Purchase orders per product line 440 340
Direct labor hours per product line 790 1,690

(a)

Correct answer iconYour answer is correct.

The total estimated manufacturing overhead was $260,000. Under traditional costing (which assigns overhead on the basis of direct labor hours), what amount of manufacturing overhead costs are assigned to: (Round answers to 2 decimal places, e.g. 12.25.)

(1) One mobile safe

$

per unit
(2) One walk-in safe

$

per unit

eTextbook and Media

Solution

Attempts: 5 of 5 used

(b1)

New attempt is in progress. Some of the new entries may impact the last attempt grading.Your answer is incorrect.

The total estimated manufacturing overhead of $260,000 was comprised of $160,000 for materials handling costs and $100,000 for purchasing activity costs. Under activity-based costing (ABC): (Round answers to 2 decimal places, e.g. 12.25.)

What amount of materials handling costs are assigned to:

(a) One mobile safe

$

per unit
(b) One walk-in safe

$

per unit
Answer #1

(a) Manufacturing overhead cost using traditional approach :-

One mobile safe = (260,000/2,480) x 790 = $82,822.58

Per unit = 82,822.58/190 = $435.91

One walk in safe = (260,000/2,480) x 1,690 = $177,177.42

Per unit = 177,177.42/40 = $4,429.44

(b1) Material handling cost :-

One mobile safe = (160,000/(190+290)) x 290 = $96,666.67

Per unit = 96,666.67/190 = $508.77

One walk in safe = (160,000/480) x 190 = $63,333.33

Per unit = 63,333.33/40 = $1,583.33